• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name E. burtin

    Log entry time 00:00:07 on October14,2005

    Entry number 156761

    keyword=evening shift summary

    16:30 5 muA CW beam in the hall for optics optimization and compton tune.
    17:50 harp scans in the arc to check beam size.
    then beam tuning in the compton until the end of the shift.
    the result of the tune is not satisfactory for the photon detector (3000 Hz/muA) and acceptable for the electron detector (15-20Hz/muA).

    MCC will continue some tests on the adiabatic damping for a while.